# Neural Operators Learn the Local Physics of Magnetohydrodynamics

### v1

@@ Abstract
Magnetohydrodynamics (MHD) plays a pivotal role in describing the dynamics of plasma and conductive fluids, essential for understanding phenomena such as the structure and evolution of stars and galaxies, and in nuclear fusion for plasma motion through ideal MHD equations. Solving these hyperbolic PDEs requires sophisticated numerical methods, presenting computational challenges due to complex structures and high costs. Recent advances introduce neural operators like the Fourier Neural Operator (FNO) as surrogate models for traditional numerical analyses. This study explores a modified Flux Fourier neural operator model to approximate the numerical flux of ideal MHD, offering a novel approach that outperforms existing neural operator models by enabling continuous inference, generalization outside sampled distributions, and faster computation compared to classical numerical schemes.
@@ 1.2 Numerical Schemes
Magnetohydrodynamics (MHD) can be viewed as a generalized model that combines hydrodynamics with electromagnetism. In particular, ideal MHD is considered as a hyperbolic conservation law, conserving mass, momentum, magnetic fields, and energy density. Numerical methods for solving hyperbolic conservation laws have a long history, with various approaches proposed to achieve stability, computational efficiency, and high-order accuracy. High-order accurate methodologies, such as essentially non-oscillatory (ENO) (Harten et al. (1987)) and weighted ENO (WENO) (Liu et al. (1994)) schemes, have been effective in solving hyperbolic conservation laws. Jiang and Shu introduced the WENO scheme with third and fifth-order accuracy (WENO-JS) (Jiang and Shu (1996)), noted for its robust shock-capturing capability, albeit with a disadvantage of being dissipative in turbulent flows. To address this shortcoming, Henrick et al. developed the enhanced WENO-M scheme (Henrick et al. (2005)). However, applying WENO schemes to the system of conservation laws involves significant computational costs. Various attempts to reduce computational costs include Jiang and Shu’s work (Jiang and Shu (1996)), Pirozzoli’s hybrid compact-WENO scheme (Pirozzoli (2002)), and the efforts by (Hill and Pullin (2004))(Costa and Don (2007)) among others. For ideal MHD, maintaining the divergence-free condition for the magnetic field is essential, with methods like the Constrained Transport (CT) method by (Evans and Hawley (1988)), and the projection scheme by (Brackbill and Barnes (1980)) proposed. Comprehensive reviews on methods ensuring the divergence-free condition in ideal MHD are available in (Toth (2000)). Early studies on numerical solutions for ideal MHD include works by (Brio and Wu (1988)), (Dai and Woodward (1998)), (Jiang and Wu (1999)). And other various efforts to solve ideal MHD numerically have been develeoped (Christlieb et al. (2014))(Rossmanith (2006)). In this research, we innovatively combine traditional numerical schemes with artificial neural network techniques to reduce computational costs and enforce the divergence-free condition for ideal MHD problems.
@@ 1.4 Our contribution
In this paper, we employ various techniques motivated from physical property of equation and numerical analysis to enhance the recently developed machine learning method, Flux FNO, for solving the ideal magnetohydrodynamic (MHD) problem. Firstly, as the problem transitions from scalar-valued to vector-valued, we redesign the Flux FNO model architecture to process each physical variable (density, velocity, magnetic field, energy) in each separated model. This is due to the experimental finding of limitations in the model’s expressiveness when handling all variables simultaneously. Secondly, we design and apply a loss function that endows the approximated numerical flux with the Total Variation Diminishing (TVD) property to ensure stability. Unlike classical numerical schemes, if we only approximate the numerical flux without such measures, severe oscillations could arise as iterations accumulate. Furthermore, we enforce the divergence-free condition by taking loss over divergence of magnetic field variables. This loss allows us to impart an inductive bias suitable for the ideal MHD solver, which we verify through an ablation study. Utilizing these methodologies, we base our model to qualitatively assess generalization performance (inference over continuous time, inference on out-of-distribution samples) and to solve representative test problems of ideal MHD, comparing them with traditional numerical methodologies.
@@ 2.1 Ideal MHD Equations
Ideal Magnetohydrodynamics (MHD) is described by a system of coupled partial differential equations that characterize conducting fluids. Among various formulations of MHD, the ideal MHD equations represent the simplest form, embodying a synthesis of fluid dynamics and Maxwell’s equations, while excluding effects such as viscosity, resistance, and thermal conductivity. The conservative form of ideal MHD can be articulated as follows:
$\rho_{t}+\nabla\cdot(\rho\mathbb{u})$ $=0$
$(\rho\mathbb{u})_{t}+\nabla\cdot\Bigg{[}\rho\mathbb{u}\otimes\mathbb{u}+\Bigg{(}p+\frac{1}{2}\|\mathbb{B}\|^{2}\Bigg{)}\mathbb{I}-\mathbb{B}\otimes\mathbb{B}\Bigg{]}$ $=0$
$\mathbb{B}_{t}+\nabla\cdot(\mathbb{u}\otimes\mathbb{B}-\mathbb{B}\otimes\mathbb{u})$ $=0$
$\mathcal{E}_{t}+\nabla\cdot\Bigg{[}\Bigg{(}\mathcal{E}+p+\frac{1}{2}\|B\|^{2}\Bigg{)}\mathbb{u}-\mathbb{B}(\mathbb{u}\cdot\mathbb{B})\Bigg{]}$ $=0$
Additionally, from Maxwell’s equations, we require the divergence-free condition for the magnetic field, which is expressed as:
$\nabla\cdot\mathbb{B}=0$
@@
$p=(\gamma-1)\bigg{(}\mathcal{E}-\frac{1}{2}\rho\mathbb{u}^{2}-\frac{1}{2}\|\mathbb{B}\|^{2}\Bigg{)}$
@@
$\frac{\partial}{\partial t}\begin{bmatrix}\rho\\ \rho u_{x}\\ \rho u_{y}\\ \rho u_{z}\\ B_{y}\\ B_{z}\\ \mathcal{E}\end{bmatrix}+\frac{\partial}{\partial x}\begin{bmatrix}\rho u_{x}\\ \rho u_{x}^{2}+p^{*}-B_{x}^{2}\\ \rho u_{x}u_{y}-B_{x}B_{y}\\ \rho u_{x}u_{z}-B_{x}B_{z}\\ B_{y}u_{x}-B_{x}u_{y}\\ B_{z}u_{x}-B_{x}u_{z}\\ (\mathcal{E}+p^{*})u_{x}-B_{x}(\mathbb{u}\cdot\mathbb{B})\end{bmatrix}=0$
@@
$\frac{\partial}{\partial t}\begin{bmatrix}\rho\\ \rho u_{x}\\ \rho u_{y}\\ \rho u_{z}\\ B_{x}\\ B_{y}\\ B_{z}\\ \mathcal{E}\end{bmatrix}+\frac{\partial}{\partial x}\begin{bmatrix}\rho u_{x}\\ \rho u_{x}^{2}+p^{*}-B_{x}^{2}\\ \rho u_{x}v_{y}-B_{x}B_{y}\\ \rho u_{x}v_{z}-B_{x}B_{z}\\ 0\\ B_{y}u_{x}-B_{x}u_{y}\\ B_{z}u_{x}-B_{x}u_{z}\\ (\mathcal{E}+p^{*})u_{x}-B_{x}(\mathbb{u}\cdot\mathbb{B})\end{bmatrix}+\frac{\partial}{\partial y}\begin{bmatrix}\rho u_{y}\\ \rho u_{y}u_{x}-B_{y}B_{x}\\ \rho u_{y}^{2}+p^{*}-B_{y}^{2}\\ \rho u_{y}u_{z}-B_{y}B_{z}\\ B_{x}u_{y}-B_{y}u_{x}\\ 0\\ B_{z}u_{y}-B_{y}u_{z}\\ (\mathcal{E}+p^{*})u_{y}-B_{y}(\mathbb{u}\cdot\mathbb{B})\end{bmatrix}=0$
@@ 2.2 Numerical schemes
$\frac{du_{j}}{dt}=-\frac{\partial f}{\partial x}\Big{|}_{x=x_{j}}$ (2)
@@
$\frac{du_{j}}{dt}=-\frac{\hat{f}_{j+\frac{1}{2}}-\hat{f}_{j-\frac{1}{2}}}{\Delta x}$ (3)
@@ 2.3 Flux Neural Operator
Among the various types of neural operators for handling functional data, the Fourier Neural Operator (FNO) is a neural operator model that processes the global convolution operations of functions by leveraging the Fourier transform. Its most notable feature is that the convolution with the kernel function is expressed through the Fourier transform as multiplication with a high-order tensor. The structure of the FNO is depicted in Figure 1, where the values of discretized functional data are non-linearly lifted by the lifting layer, passed through the Fourier layers, and finally projected into the desired dimension vector by the projection layer composed of FNCs. As shown in diagram (b) of Figure 1, each Fourier layer consists of a function convolution operation that globally transforms the data and an auxiliary neural network, which must maintain resolution invariance, with CNN layers commonly adopted to handle local data processing. The mathematical formulation for a Fourier Neural Operator (FNO) with a CNN layers can be represented as follows:
@@
$v_{0}$ $:=\mathcal{N}_{P}(a|_{X})=(\mathcal{N}_{P}(a_{\textbf{x}\cdot})_{j})_{\textbf{x}\in\textbf{X},j=1,\dots,d_{v_{0}}}$
@@
$+\sum_{\textbf{z},\textbf{k}\in K,k}{D^{{\dagger}}_{\textbf{x}\textbf{k}}R_{t+1,\textbf{k},jk}D_{\textbf{k}\textbf{z}}v_{t,\textbf{z}k}}\Big{)}\quad(t=0,...,L-1)$
$G(a;\theta)$ $:=\mathcal{N}_{Q}(v_{L})=(\mathcal{N}_{Q}(v_{L\textbf{x}\cdot})_{j})_{\textbf{x}\in\textbf{X},j=1,\dots,d_{v_{L}}}$
where $\mathcal{N}_{P}$ and $\mathcal{N}_{Q}$ are neural networks used for lifting and projection, respectively. $a|_{X}$ represents the discretized functional data of $a$, and each $C_{i}$ is a $d$-dimensional CNN layer with a kernel tensor $K_{i}$. The $R_{i}$ tensors are weight tensors that parameterize the kernel function of the Fourier layers. $D_{\textbf{k}\textbf{z}}$ represents the components of the discrete Fourier transform, also denoted as $\mathcal{F}$.
@@
$\mathbb{U}_{j+1}=\mathbb{U}_{j}+\sum_{i=1}^{k}\frac{\Delta t}{\Delta x_{i}}\Bigg{[}G_{i}(\mathbb{U}_{j}^{l_{i,1}},\dots,\mathbb{U}_{j}^{l_{i,m}};\theta_{i})-G_{i}(\mathbb{U}_{j}^{r_{i,1}},\dots,\mathbb{U}_{j}^{r_{i,m}};\theta_{i})\Bigg{]}$ (6)
@@
$-\Bigg{(}\sum_{i=1}^{k}\frac{\Delta t}{\Delta x_{i}}\Bigg{[}G_{i}(\mathbb{U}_{i,j}^{l_{i,1}},\dots,\mathbb{U}_{i,j}^{l_{i,m}};\theta_{i})-G_{i}(\mathbb{U}_{i,j}^{r_{i,1}},\dots,\mathbb{U}_{i,j}^{r_{i,m}};\theta_{i})\Bigg{]}\Bigg{)}\Bigg{\|}^{2}$ (7)
@@
$\mathcal{L}_{consi}(\{\mathbb{U}_{i}\}_{j=1}^{B}):=\sum_{i=1}^{B}\sum_{j=1}^{N^{t}}\sum_{l=1}^{k}\|G_{l}(\mathbb{U}_{i,j},\dots,\mathbb{U}_{i,j};\theta_{l})-F_{l}(\mathbb{U}_{i,j})\|^{2}$ (8)
@@ 3.1 Adaptive FNO Architecture for Multidimensional Outputs in Ideal MHD
In the referenced study (Kim and Kang (2024a)), the output dimension was one-dimensional since it dealt with one-dimensional scalar conservation laws. However, in the case of ideal Magnetohydrodynamics (MHD) discussed in this paper, the output dimension is seven-dimensional for one-dimensional problems and eight-dimensional for two-dimensional problems. We experimentally observed significant issues with expressiveness when a single FNO model was tasked with handling all outputs in cases of large output dimensions. To address this issue, we allocated separate FNO models to handle each physical quantity. Specifically, density, energy, velocity vectors, and magnetic field vectors were each processed by their own FNO models. For two-dimensional problems, given that the physical flux functions exist for both the x and y axes, we further segregated the considerations for each physical flux. This architectural approach is summarized in the schematic shown in the Figure 3. Modified version of (6) for one-dimensional case can be written as follows:
@@
$G_{\mathbb{B}}(\mathbb{U}_{j}^{r_{1}},\dots,\mathbb{U}_{j}^{r_{m}};\theta_{\mathbb{B}})^{T},G_{E}(\mathbb{U}_{j}^{r_{1}},\dots,\mathbb{U}_{j}^{r_{m}};\theta_{E})\Big{)}^{T}$
@@
$F_{\mathbb{B}}(\mathbb{U}_{j}^{b_{1}},\dots,\mathbb{U}_{j}^{b_{m}};\theta_{\mathbb{B}}^{F})^{T},F_{E}(\mathbb{U}_{j}^{b_{1}},\dots,\mathbb{U}_{j}^{b_{m}};\theta_{E}^{F})\Big{)}^{T}$
@@
$\mathcal{L}_{tm}(\{\mathbb{U}_{i}\}_{i=1}^{B})$ $:=\sum_{i=1}^{B}\sum_{j=1}^{N_{t}-1}\Big{\|}(\mathbb{U}_{i,j+1}-\mathbb{U}_{i,j})-\Delta\mathbb{R}(\mathbb{U}_{i,j};\theta)\Big{\|}^{2}$ (11)

### v2

@@ Abstract
Magnetohydrodynamics (MHD) plays a pivotal role in describing the dynamics of plasma and conductive fluids, essential for understanding phenomena such as the structure and evolution of stars and galaxies, and in nuclear fusion for plasma motion through ideal MHD equations. Solving these hyperbolic PDEs requires sophisticated numerical methods, presenting computational challenges due to complex structures and high costs. Recent advances introduce neural operators like the Fourier Neural Operator (FNO) as surrogate models for traditional numerical analyses. This study explores a modified Flux Neural Operator (Flux NO) model to approximate the numerical flux of ideal MHD, offering a novel approach that outperforms existing neural operator models by enabling continuous inference, generalization outside sampled distributions, and faster computation compared to classical numerical schemes.
Keywords: Scientific machine learning, Neural operator, Numerical analysis, Magnetohydrodynamics
@@ 1.2 Numerical Schemes
MHD can be viewed as a generalized model that combines hydrodynamics with electromagnetism. In particular, ideal MHD is considered as a hyperbolic conservation law, conserving mass, momentum, magnetic fields, and energy density. Numerical methods for solving hyperbolic conservation laws have a long history, with various approaches proposed to achieve stability, computational efficiency, and high-order accuracy. High-order accurate methodologies, such as essentially non-oscillatory (ENO) (Harten et al. (1987)) and weighted ENO (WENO) (Liu et al. (1994)) schemes, have been effective in solving hyperbolic conservation laws. Jiang and Shu introduced the WENO scheme with third and fifth-order accuracy (WENO-JS) (Jiang and Shu (1996)), noted for its robust shock-capturing capability, albeit with a disadvantage of being dissipative in turbulent flows. To address this shortcoming, Henrick et al. developed the enhanced WENO-M scheme (Henrick et al. (2005)). However, applying WENO schemes to the system of conservation laws involves significant computational costs. Various attempts to reduce computational costs include Jiang and Shu’s work (Jiang and Shu (1996)), Pirozzoli’s hybrid compact-WENO scheme (Pirozzoli (2002)), and the efforts by (Hill and Pullin (2004))(Costa and Don (2007)) among others. For ideal MHD, maintaining the divergence-free condition for the magnetic field is essential, with methods like the Constrained Transport (CT) method by (Evans and Hawley (1988)), and the projection scheme by (Brackbill and Barnes (1980)) proposed. Comprehensive reviews on methods ensuring the divergence-free condition in ideal MHD are available in (Toth (2000)). Early studies on numerical solutions for ideal MHD include works by (Brio and Wu (1988)), (Dai and Woodward (1998)), (Jiang and Wu (1999)). And other various efforts to solve ideal MHD numerically have been develeoped (Christlieb et al. (2014))(Rossmanith (2006)). In this research, we innovatively combine traditional numerical schemes with artificial neural network techniques to reduce computational costs and enforce the divergence-free condition for ideal MHD problems.
@@ 1.4 Our contribution
In this paper, we employ various techniques motivated by the physical properties of equations and numerical analysis to enhance the recently developed machine learning method, Flux FNO, for solving the ideal MHD problem. As the problem transitions from scalar-valued to vector-valued outputs, we have redesigned the Flux FNO model architecture. This redesign allows the model to process each physical variable—density, velocity, magnetic field, energy—separately, addressing the experimental findings that revealed limitations in the model’s expressiveness when handling all variables simultaneously. Secondly, we design and apply a loss function that confers the Total Variation Diminishing (TVD) property to the approximated numerical flux, ensuring stability. Unlike classical numerical schemes, merely approximating the numerical flux without such measures could lead to severe oscillations as iterations accumulate. Furthermore, we enforce the divergence-free condition by implementing a loss specific to the divergence of magnetic field variables. This specialized loss imparts an inductive bias suitable for the ideal MHD solver, effectiveness of which we confirm through an ablation study. Utilizing these methodologies, our experiments are designed to qualitatively assess its generalization performance, including inference over continuous time and on out-of-distribution samples. We also apply it to solve representative test problems of ideal MHD, comparing its effectiveness with traditional numerical methodologies.
@@ 2.1 Ideal MHD Equations
Ideal MHD is described by a system of coupled partial differential equations that characterize conducting fluids. Among various formulations of MHD, the ideal MHD equations represent the simplest form, embodying a synthesis of fluid dynamics and Maxwell’s equations, while excluding effects such as viscosity, resistance, and thermal conductivity. The conservative form of ideal MHD can be articulated as follows:
$\rho_{t}+\nabla\cdot(\rho\mathbb{u})$ $=0,$
$(\rho\mathbb{u})_{t}+\nabla\cdot\Bigg{[}\rho\mathbb{u}\otimes\mathbb{u}+\Bigg{(}p+\frac{1}{2}\|\mathbb{B}\|^{2}\Bigg{)}\mathbb{I}-\mathbb{B}\otimes\mathbb{B}\Bigg{]}$ $=0,$
$\mathbb{B}_{t}+\nabla\cdot(\mathbb{u}\otimes\mathbb{B}-\mathbb{B}\otimes\mathbb{u})$ $=0,$
$\mathcal{E}_{t}+\nabla\cdot\Bigg{[}\Bigg{(}\mathcal{E}+p+\frac{1}{2}\|B\|^{2}\Bigg{)}\mathbb{u}-\mathbb{B}(\mathbb{u}\cdot\mathbb{B})\Bigg{]}$ $=0.$
where $\mathbb{u}$ represents velocity, $\mathbb{B}$ the magnetic field, $\rho$ density, $E$ energy, and $p$ pressure. Additionally, according to Maxwell’s equations, the magnetic field must satisfy the divergence-free condition, which is expressed as:
$\nabla\cdot\mathbb{B}=0.$
@@
$p=(\gamma-1)\bigg{(}\mathcal{E}-\frac{1}{2}\rho\mathbb{u}^{2}-\frac{1}{2}\|\mathbb{B}\|^{2}\Bigg{)}.$
@@
$\frac{\partial}{\partial t}\begin{bmatrix}\rho\\ \rho u_{x}\\ \rho u_{y}\\ \rho u_{z}\\ B_{y}\\ B_{z}\\ \mathcal{E}\end{bmatrix}+\frac{\partial}{\partial x}\begin{bmatrix}\rho u_{x}\\ \rho u_{x}^{2}+p^{*}-B_{x}^{2}\\ \rho u_{x}u_{y}-B_{x}B_{y}\\ \rho u_{x}u_{z}-B_{x}B_{z}\\ B_{y}u_{x}-B_{x}u_{y}\\ B_{z}u_{x}-B_{x}u_{z}\\ (\mathcal{E}+p^{*})u_{x}-B_{x}(\mathbb{u}\cdot\mathbb{B})\end{bmatrix}=0.$
@@
$\frac{\partial}{\partial t}\begin{bmatrix}\rho\\ \rho u_{x}\\ \rho u_{y}\\ \rho u_{z}\\ B_{x}\\ B_{y}\\ B_{z}\\ \mathcal{E}\end{bmatrix}+\frac{\partial}{\partial x}\begin{bmatrix}\rho u_{x}\\ \rho u_{x}^{2}+p^{*}-B_{x}^{2}\\ \rho u_{x}v_{y}-B_{x}B_{y}\\ \rho u_{x}v_{z}-B_{x}B_{z}\\ 0\\ B_{y}u_{x}-B_{x}u_{y}\\ B_{z}u_{x}-B_{x}u_{z}\\ (\mathcal{E}+p^{*})u_{x}-B_{x}(\mathbb{u}\cdot\mathbb{B})\end{bmatrix}+\frac{\partial}{\partial y}\begin{bmatrix}\rho u_{y}\\ \rho u_{y}u_{x}-B_{y}B_{x}\\ \rho u_{y}^{2}+p^{*}-B_{y}^{2}\\ \rho u_{y}u_{z}-B_{y}B_{z}\\ B_{x}u_{y}-B_{y}u_{x}\\ 0\\ B_{z}u_{y}-B_{y}u_{z}\\ (\mathcal{E}+p^{*})u_{y}-B_{y}(\mathbb{u}\cdot\mathbb{B})\end{bmatrix}=0.$
@@ 2.2 Numerical schemes
$\frac{du_{j}}{dt}=-\frac{\partial f}{\partial x}\Big{|}_{x=x_{j}}.$ (2)
@@
$\frac{du_{j}}{dt}=-\frac{\hat{f}_{j+\frac{1}{2}}-\hat{f}_{j-\frac{1}{2}}}{\Delta x}.$ (3)
@@ 2.3 Flux Neural Operator
Among the various types of neural operators for handling functional data, the FNO is a neural operator model that processes the global convolution operations of functions by leveraging the Fourier transform. Its most notable feature is that the convolution with the kernel function is expressed through the Fourier transform as multiplication with a high-order tensor. The structure of the FNO is depicted in Figure 1, where the values of discretized functional data are non-linearly lifted by the lifting layer, passed through the Fourier layers, and finally projected into the desired dimension vector by the projection layer composed of FNCs. As shown in diagram (b) of Figure 1, each Fourier layer consists of a function convolution operation that globally transforms the data and an auxiliary neural network, which must maintain resolution invariance, with CNN layers commonly adopted to handle local data processing. The mathematical formulation for a FNO with a CNN layers can be represented as follows:
@@
$v_{0}$ $:=\mathcal{N}_{P}(a|_{X})=(\mathcal{N}_{P}(a_{\textbf{x}\cdot})_{j})_{\textbf{x}\in\textbf{X},j=1,\dots,d_{v_{0}}},$
@@
$+\sum_{\textbf{z},\textbf{k}\in K,k}{D^{{\dagger}}_{\textbf{x}\textbf{k}}R_{t+1,\textbf{k},jk}D_{\textbf{k}\textbf{z}}v_{t,\textbf{z}k}}\Big{)},\quad(t=0,...,L-1)$
$G(a;\theta)$ $:=\mathcal{N}_{Q}(v_{L})=(\mathcal{N}_{Q}(v_{L\textbf{x}\cdot})_{j})_{\textbf{x}\in\textbf{X},j=1,\dots,d_{v_{L}}}.$
Here, $\mathcal{N}_{P}$ and $\mathcal{N}_{Q}$ are neural networks used for lifting and projection, respectively, while $\mathcal{A}_{t+1}$ represents Fourier layers with $L$ being the depth of these Fourier layers. $a|_{X}$ represents the discretized functional data of $a$, $\sigma$ is an activation function, and each $C_{i}$ is a $d$-dimensional CNN layer with a kernel tensor $K_{i}$. The $R_{i}$ tensors are weight tensors that parameterize the kernel function of the Fourier layers. $D_{\textbf{k}\textbf{z}}$ represents the components of the discrete Fourier transform, also denoted as $\mathcal{F}$.
@@
$\mathbb{U}_{j+1}=\mathbb{U}_{j}+\sum_{i=1}^{k}\frac{\Delta t}{\Delta x_{i}}\Bigg{[}G_{i}(\mathbb{U}_{j}^{l_{i,1}},\dots,\mathbb{U}_{j}^{l_{i,m}};\theta_{i})-G_{i}(\mathbb{U}_{j}^{r_{i,1}},\dots,\mathbb{U}_{j}^{r_{i,m}};\theta_{i})\Bigg{]}.$ (6)
@@
$-\Bigg{(}\sum_{i=1}^{k}\frac{\Delta t}{\Delta x_{i}}\Bigg{[}G_{i}(\mathbb{U}_{i,j}^{l_{i,1}},\dots,\mathbb{U}_{i,j}^{l_{i,m}};\theta_{i})-G_{i}(\mathbb{U}_{i,j}^{r_{i,1}},\dots,\mathbb{U}_{i,j}^{r_{i,m}};\theta_{i})\Bigg{]}\Bigg{)}\Bigg{\|}^{2}.$ (7)
@@
$\mathcal{L}_{consi}(\{\mathbb{U}_{i}\}_{j=1}^{B}):=\sum_{i=1}^{B}\sum_{j=1}^{N^{t}}\sum_{l=1}^{k}\|G_{l}(\mathbb{U}_{i,j},\dots,\mathbb{U}_{i,j};\theta_{l})-F_{l}(\mathbb{U}_{i,j})\|^{2}.$ (8)
@@ 3.1 Adaptive FNO Architecture for Multidimensional Outputs in Ideal MHD
In the referenced study (Kim and Kang (2024a)), the output dimension was one-dimensional since it dealt with one-dimensional scalar conservation laws. However, in the case of ideal MHD discussed in this paper, the output dimension is seven-dimensional for one-dimensional problems and eight-dimensional for two-dimensional problems. We experimentally observed performance degradation when a single FNO model was tasked with handling all outputs in cases of large output dimensions; the results of these experiments can be found in Table 7. To address this issue, we allocated separate FNO models to handle each physical quantity. Specifically, density, energy, velocity vectors, and magnetic field vectors were each processed by their own FNO models. For two-dimensional problems, given that the physical flux functions exist for both the x and y axes, we further segregated the considerations for each physical flux. This architectural approach is summarized in the schematic shown in the Figure 3. Modified version of (6) for one-dimensional case can be written as follows:
@@
$G_{\mathbb{B}}(\mathbb{U}_{j}^{r_{1}},\dots,\mathbb{U}_{j}^{r_{m}};\theta_{\mathbb{B}})^{T},G_{E}(\mathbb{U}_{j}^{r_{1}},\dots,\mathbb{U}_{j}^{r_{m}};\theta_{E})\Big{)}^{T}.$
@@
$F_{\mathbb{B}}(\mathbb{U}_{j}^{b_{1}},\dots,\mathbb{U}_{j}^{b_{m}};\theta_{\mathbb{B}}^{F})^{T},F_{E}(\mathbb{U}_{j}^{b_{1}},\dots,\mathbb{U}_{j}^{b_{m}};\theta_{E}^{F})\Big{)}^{T}.$
@@
$\mathcal{L}_{tm}(\{\mathbb{U}_{i}\}_{i=1}^{B})$ $:=\sum_{i=1}^{B}\sum_{j=1}^{N_{t}-1}\Big{\|}(\mathbb{U}_{i,j+1}-\mathbb{U}_{i,j})-\Delta\mathbb{R}(\mathbb{U}_{i,j};\theta)\Big{\|}^{2}.$ (11)